OpenStudy (anonymous):

AWARDED MEDAL FOR HELP/FANNED You deposit $1,000 in an account at the Lifelong Trust Savings and Loan that pays 5% interest compounded quarterly. By how much will your deposit have grown after 9 years? Round your answer to the nearest cent.

OpenStudy (wolf1728):

Total = Principal * (1 + (rate/n)^n*t where 'n' is the number of compounding periods Total = 1,000 * (1+.05/4)4*9 Total = 1,000 * 1.0125^36 Total = 1,000 * 1.5639438187 Total = 1,563.94 You may check your answer here: http://www.1728.org/compint.htm

OpenStudy (anonymous):

Thank you!

OpenStudy (wolf1728):

u r welcome